OpenStudy (anonymous):

A party rental company has chairs and tables for rent. The total cost to rent 5 chairs and 3 tables is $31 . The total cost to rent 2 chairs and 6 tables is $52 . What is the cost to rent each chair and each table?

OpenStudy (anonymous):

the chairs could be a dollar and the tables are 8.333

OpenStudy (gudden):

let the cost of renting a chair be x and for renting a table be y than 5x + 3y = 31 and 2x + 6y = 52 2x+ 6y = 52 can be further simplified to x + 3y = 26 now try solving for x and y... best of luck:D